OpenStudy (anonymous):

If x= y-3 and y=z^2, what is x in terms of z?

hartnn (hartnn):

just plug in z^2 in place of y so, x=z^2 -3 thats it.

OpenStudy (hba):

Just plug in y=z^2 in x=y-3